A substance of density 4 grams per centimeter cube(g/cm3) causes 60g of turpentine (density 0.80 grams per centimeter cube(g/cm3)) to overflow from a displacement vessel. What is the mass of the substance?
1
Expert's answer
2017-01-11T12:41:10-0500
ρ=m/V V=m/ρ m = ρ·V
V(turpentine displaced) = 60/0.80= 75 cm3
m (solid) = 4 · 75 = 300 g